# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@ 2.3 $h^{*}$- and local $h^{*}$-polynomials of lattice polytopes
where $h^{*}(P;t)$ is a polynomial of degree at most $\dim P$ called the $\mathbf{h^{*}}$-polynomial of $P$ (see, for example, [BR07, Section 3.3]). A non-trivial theorem of Stanley [Stan80] states that the coefficients of the $h^{*}$-polynomial are non-negative integers. Throughout, we will use the notation
@@
(see e.g. [KS16, Example 7.1]) The $h^{*}$-polynomial of lattice polytope $P\subset\mathbb{Z}^{n}$ satisfy the following properties:
@@
The following definition was introduced by Stanley in [Stan92, Example 7.13], generalizing the definition of Betke and McMullen in the case of a simplex [BeM85]. It was independently introduced by Borisov and Mavlyutov in [BoM03] as $\widetilde{S}(t)$.
@@
[KS16, Lemma 7.3(2)] Let $P$ be a non-empty lattice polytope. Then the $h^{*}$-polynomial of a polytope can be recovered from the local $h^{*}$-polynomials of its faces via
@@
Notation 2.27.
@@
In [KS16] and some other papers it is called the ‘‘link’’ but the ‘‘star’’ is a more common way to name it.
@@
The local $h^{*}$-polynomial does not have unimodular coefficients in general (see e.g. [KS16, Example 7.22.]). However it has unimodular coefficients for IDP polytopes [APPS22, Theorem 2.5.] and $s$-lecture hall polytopes [GS20, Corollary 5.5.].
@@ 3 Strong formal subdivision with boundary
In this section we define strong formal subdivisions with boundary, which are strong formal subdivision with marked elements, called boundary elements. The restriction to these boundary elements must also be a strong formal subdivision.
We define the agglutination of the strong formal subdivisions with the same boundaries along their boundary. Lemma 3.8 provides a simple formula for the local $h$-polynomial of the agglutination in terms of the local $h$-polynomials of the initial strong formal subdivisions and their boundary. The most important examples (Example 3.9) of strong formal subdivision with boundary are obtained from polyhedral subdivisions of convenient polytopes inside a cone. We prove that the local $h$-polynomial of such agglutinations inherit the properties of the local $h$-polynomial of usual polyhedral subdivisions, including nonnegativity and symmetricity.
@@ 3.1 Poset with boundary and twins-poset
In this subsection we define two types of posets and prove their basic properties. We need them to define strong formal subdivisions with boundaries. Recall that a lower set of a poset is a subposet which, together with each element, contains all smaller elements from the original poset.
@@
Consider two posets $(\Gamma^{1},\Gamma_{B})$ and $(\Gamma^{2},\Gamma_{B})$ with common boundary $\Gamma_{B}=\Gamma^{1}\cap\Gamma^{2}$. The agglutination $\Gamma^{1}\sharp_{\Gamma_{B}}\Gamma^{2}$ is the poset $\Gamma^{1}\cup\Gamma^{2}$ with incomparable $\Gamma^{1}\setminus\Gamma_{B}$ and $\Gamma^{2}\setminus\Gamma_{B}$, and the induced poset structure on both $\Gamma^{1}$ and $\Gamma^{2}$.
Note that the agglutination of any pair of lower Eulerian posets is lower Eulerian.
@@
$h(\Gamma^{1}\sharp_{\Gamma_{B}}\Gamma^{2};t)=h(\Gamma^{1};t)+h(\Gamma^{2};t)+(t-1)h(\Gamma_{B};t)$
@@
$h(\Gamma^{1}\sharp_{\Gamma_{B}}\Gamma^{2};t)$ $=(h(\Gamma^{1};t)-t(t^{-1}-1)h(\Gamma_{B};t))+(h(\Gamma^{2};t)-t(t^{-1}-1)h(\Gamma_{B};t))+$
@@
By Claim 2.8, we have:
$t^{n-1}h(R\setminus\hat{1}_{R};t^{-1})=h(R\setminus\hat{1}_{R};t)=\frac{g(R;t)-t^{n}g(R;t)}{1-t},$
@@
This implies that both sides vanish, and thus $g(R^{\pm},t)=g(R;t)$.
@@ 3.2 Strong formal subdivisions with boundaries
Note that the restriction of $\sigma$ on $\Gamma\setminus\Gamma_{B}$ is generally not a strong formal subdivision since $\Gamma\setminus\Gamma_{B}$ may not be lower Eulerian (may have several minimal elements).
@@
Applying Lemma 3.5, we obtain that Lemma 3.8 is implied by the following equation. To prove the lemma, it suffices to show that for every $x\in R$ we have the equality:
@@ 3.3 Subdivisions of convenient polytopes in cones
For a cone $C$, denote its face poset by $\mathcal{C}$ as in Example 2.9. Note that it is an Eulerian poset.
@@
Note that the polytope’s convenience implies that the map $\sigma$ is strongly surjective. The remaining conditions defining strong formal subdivisions can be confirmed the same way as for polyhedral subdivisions of polytopes [KS16, Lemma 3.25.].
Consider the face poset $\mathcal{P}$ of polytope $P$.
@@
Note that $\mathcal{P}_{B}$ can be obtained from regular polyhedral subdivision:
@@
    Let $C_{\min}$ the minimal non-empty face of $C$ (i.e., its maximal affine subspace).
@@
    Define the height function: 00 on the vertices of $P$, and $1$ at $X$.
@@
Then for any $F\in\mathcal{P}_{B}$ we have: $\operatorname{St}_{\mathcal{P}_{B}}(F)=\operatorname{St}_{\mathcal{X}}(\operatorname{Conv}(F,X))$. Thus, for any face $Q\in\mathcal{C}:F\subset Q$ we have $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,F;t)=\ell_{\mathcal{P}}(\mathcal{X},Q,F;t)$.
Consider a face $U\in\mathcal{P}_{B}$ of $P$ and face $Q\in\mathcal{C}:U\subset Q$. Denote $\dim(U)=u$ and $\dim(Q)=q$.
@@
where $\deg(R_{<(q-u)/2})<(q-u)/2$ is completely determined by $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$.
@@
The polynomial $\ell_{\mathcal{C}^{\pm}}(\mathcal{P},Q,U;t)$ has symmetric coefficients centered at $(q-u)/2$ and polynomial $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$ has symmetric coefficients centered at $(q-u-1)/2$. So Lemma 3.11 fully recovers $\ell_{\mathcal{C}^{\pm}}(\mathcal{P},Q,U;t)$ from $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$. For instance, $\ell_{\mathcal{C}^{\pm}}(\mathcal{P},Q,U;t)$ has non-positive coefficients.
@@
    The first summand is the the sum of products of pairs of $g$-polynomials of posets with rank sum $q-u+1$.
@@
Thus, the first summand has degree less than $(q-u)/2$ and can be rewritten as $R_{<(q-u)/2}$.
The second summand is equal to $-\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$ and this completes the proof.
@@
Recall that by Corollary 2.16 the local $h$-polynomial of $\ell_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P},Q,U;t)$ has symmetric coefficients centered at $(q-u)/2$, so they are indeed determined by $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$.
By Theorem 2.22 and Remark 3.10 the polynomial $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$ has symmetric unimodal coefficients centered at $(q-u-1)/2$. Hence, we obtained the following corollary.
@@
4 Generalization of $h^{*}$- and $\ell^{*}$-polynomial
4.1 $h^{*}$- and $\ell^{*}$-polynomial of polyhedral strong formal subdivision
We call an abstract lattice polyhedral complex $\mathcal{S}$ a poset of ‘‘abstract’’ convex lattice polytopes, such that every face of any polytope is also contained in the poset and any two polytopes intersect along a common face. Formally, abstract lattice polyhedral complex is a map $\theta$ from a poset $\Gamma$ to lattice polytopes in different lattices such that for every $x\in\Gamma$ the lower set $\Gamma_{\leq x}$ is the face poset of the polytope $\theta(x)$ and for every $y\leq x$ there is a lattice isomorphism between the polytope $\theta(y)$ and the corresponding face of the polytope $\theta(x)$. We do not assume that there is a geometric realization in a common lattice.
@@ 4.1 $h^{*}$- and $\ell^{*}$-polynomial of polyhedral strong formal subdivision
It follows e.g. from [KS16, Theorem 3.11] that we recover the $h^{*}$-polynomial via the local $h^{*}$-polynomial
@@
The first formula is implied by Definition 4.1 and the corresponding Katz-Stapledon decomposition formula from Lemma 2.29. The equivalence of these formulas is proved the same way as in [KS16, Lemma 7.12.]. ∎
4.2 $h^{*}$- and $\ell^{*}$-polynomial of agglutination

### v2

@@ 2.3 $\mathbf{h^{*}}$- and local $\mathbf{h^{*}}$-polynomials of lattice polytopes
where $h^{*}(P;t)$ is a polynomial of degree at most $\dim P$ called the $\mathbf{h^{*}}$-polynomial of $P$ (see, for example, [BR07, Section 3.4]). A non-trivial theorem of Stanley [Stan80, Theorem 2.1] states that the coefficients of the $h^{*}$-polynomial are non-negative integers. Throughout, we will use the notation
@@
(see, e.g., [KS16, Example 7.1]) The $h^{*}$-polynomial of lattice polytope $P\subset\mathbb{Z}^{n}$ satisfy the following properties:
@@
The following definition was introduced by Stanley in [Stan92, Example 7.13], generalizing the definition of Betke and McMullen in the case of a simplex [BeM85]. It was independently introduced by Borisov and Mavlyutov in [BoM03] as $\widetilde{S}(t)$ in the context of the Batyrev’s construction of mirror symmetry [Bat94].
@@
[KS16, see, e.g., Lemma 7.3(2)] Let $P$ be a non-empty lattice polytope. Then the $h^{*}$-polynomial of a polytope can be recovered from the local $h^{*}$-polynomials of its faces via
@@
The $h^{*}$- and local $h^{*}$-polynomials can be considered as the explicit output of the algorithm for computing the $\chi_{y}$-characteristic and the Hodge-Deligne numbers from [DK86]. For more details see the remarks below.
Remark 2.27.
In [DK86] the notations of $l$ and $l^{*}$ have nothing in common with the local $h$- and $h^{*}$-polynomials (which were defined later in [Stan92]). In [DK86] these notations just count lattice points in the polytope and in its interior.
Remark 2.28.
(There is a typo in the formula from [KS16a, Theorem 5.5].) Consider polynomial $f$ with Newton polytope $P$ and generic coefficients. Then the $h^{*}$-polynomial of $P$ computes the $\chi_{y}$-characteristic of $\{f=0\}\subset(\mathbb{C}\setminus 0)^{n}$ via:
$y\chi_{y}(\{f=0\})=yE(\{f=0\};-y,1)=(-1)^{\dim P+1}(y+1)^{\dim P}+(-1)^{\dim P}h^{*}(P;-y).$
Remark 2.29.
Define the mixed $h^{*}$-polynomial via the formula (from [KS16, Definition 7.5]):
$h^{*}(P;u,v)=\sum\limits_{Q\subseteq P}v^{\dim Q+1}\ell^{*}(Q,uv^{-1})g([Q,P];uv)$
Consider polynomial $f$ with Newton polytope $P$ and generic coefficients. Then (see [KS16a, §5.4]) we have the following formula for the Hodge-Deligne Polynomial (or E-polynomial) of the hypersurface $\{f=0\}\subset(\mathbb{C}\setminus 0)^{n}$:
$uvE(\{f=0\};u,v)=(uv-1)^{\dim P}+(-1)^{\dim P+1}h^{*}(P;u,v).$
For more general formulas (including limit mixed Hodge structures) see [KS16a].
Remark 2.30.
Similar combinatorics explicitly computes ‘‘stringy invariant’’ in the Batyrev’s construction of mirror symmetry [Bat94]. This invariant was computed by Batyrev and Borisov in [BB96, Theorem 4.14] and simplified by Borisov and Mavlyutov in [BoM03, Theorem 7.2]. For more details, see [KS16a, Example 5.13].
Notation 2.31.
@@
In [KS16] and [BKN24] it is called the ‘‘link’’ but the ‘‘star’’ is a more common way to name it.
@@
The local $h^{*}$-polynomial does not have unimodular coefficients in general (see e.g. [KS16, Example 7.22]). However it has unimodular coefficients for IDP polytopes [APPS22, Theorem 2.5] and $s$-lecture hall polytopes [GS20, Corollary 5.5].
@@ 3 Strong formal subdivision with boundary
In this section we define strong formal subdivisions with boundary. These are strong formal subdivisions with a specified lower set of elements, called the boundary. The restriction of the subdivision to these boundary elements must itself be a strong formal subdivision.
We define the agglutination of the strong formal subdivisions with the same boundaries along their boundary. Lemma 3.8 provides a simple formula for the local $h$-polynomial of the agglutination in terms of the local $h$-polynomials of the initial strong formal subdivisions and their boundary. The most important examples (Example 3.9) of strong formal subdivision with boundary are obtained from polyhedral subdivisions of convenient polytopes inside a cone. We prove that the local $h$-polynomial of such agglutinations inherits the properties of the local $h$-polynomial of usual polyhedral subdivisions, including nonnegativity and symmetricity.
@@ 3.1 Poset with boundary and twins-poset
In this subsection we define two types of posets and prove their basic properties. We need them to define strong formal subdivisions with boundaries. Recall that a lower set of a poset is a subposet which, together with each element, contains all smaller elements.
@@
Consider two posets $(\Gamma^{1},\Gamma_{B})$ and $(\Gamma^{2},\Gamma_{B})$ with common boundary $\Gamma_{B}=\Gamma^{1}\cap\Gamma^{2}$. The agglutination $\Gamma^{1}\sharp_{\Gamma_{B}}\sharp\Gamma^{2}$ is the poset $\Gamma^{1}\cup\Gamma^{2}$ with incomparable $\Gamma^{1}\setminus\Gamma_{B}$ and $\Gamma^{2}\setminus\Gamma_{B}$, and the induced poset structure on both $\Gamma^{1}$ and $\Gamma^{2}$.
The agglutination of two (lower) Eulerian posets with a common boundary is itself (lower) Eulerian (this follows from verifying the defining property on intervals).
@@
$h(\Gamma^{1}\sharp_{\Gamma_{B}}\sharp\Gamma^{2};t)=h(\Gamma^{1};t)+h(\Gamma^{2};t)+(t-1)h(\Gamma_{B};t)$
@@
$h(\Gamma^{1}\sharp_{\Gamma_{B}}\sharp\Gamma^{2};t)$ $=(h(\Gamma^{1};t)-t(t^{-1}-1)h(\Gamma_{B};t))+(h(\Gamma^{2};t)-t(t^{-1}-1)h(\Gamma_{B};t))+$
@@
Figure 1: Hasse diagram of a poset (left) and the twins-poset (right) born from it
@@
By Remark 2.8, we have:
$t^{n-1}h(R\setminus\hat{1}_{R};t^{-1})=h(R\setminus\hat{1}_{R};t)=\frac{g(R;t)-t^{n}g(R;t^{-1})}{1-t},$
@@
The left-hand side is a polynomial whose terms have degree greater than $(n+1)/2$, while the right-hand side is a polynomial whose terms have degree less than $(n+1)/2$. The only polynomial that can satisfy this identity is the zero polynomial. Therefore, both sides must be zero, which implies $g(R^{\pm},t)=g(R;t)$..
@@ 3.2 Strong formal subdivisions with boundaries
Our main examples strong formal subdivision with boundaries come from convenient polytopes in cones (see Example 3.9).
Note that the restriction of $\sigma$ to $\Gamma\setminus\Gamma_{B}$ is generally not a strong formal subdivision, since $\Gamma\setminus\Gamma_{B}$ may not be lower Eulerian (it can have multiple minimal elements).
@@
To prove the lemma, it suffices to show that for every $x\in R$ we have the equality:
@@
Indeed, if this holds, then substituting into the definition of $\ell_{R}(\Gamma^{1}\sharp_{\Gamma_{B}}\sharp\Gamma^{2};t)$ and applying Lemma 3.5 would yield the desired formula.
@@ 3.3 Subdivisions of convenient polytopes in cones
For a cone $C$, denote its Eulerian face poset by $\mathcal{C}$ as in Example 2.9.
@@
Figure 2: Convenient square $P=OABC$ in the cone $C=\mathbb{R}^{2}_{\geq 0}$ with red boundary $\partial_{C}(P)$
Figure 3: Strong formal subdivision (dashed vectors) between the face poset $\mathcal{P}$ of the square OABCOABC (right) and $\mathcal{C}^{\pm}$ (left). Hasse diagrams of the boundary poset $\mathcal{P}_{B}\subset\mathcal{P}$, the lower set $\mathcal{C}^{-}\subset\mathcal{C}^{\pm}$ and the mapping between them are drawn red.
Note that the polytope’s convenience implies that the map $\sigma$ is strongly surjective. The remaining conditions defining strong formal subdivisions can be confirmed the same way as for polyhedral subdivisions of polytopes [KS16, Lemma 3.25].
@@
Consider the face poset $\mathcal{P}$ of a convenient polytope $P$ in a cone $C$. Note that $\mathcal{P}_{B}$ (the boundary poset from Example 3.9 for $\mathcal{S}=\mathcal{P}$) can be obtained from regular polyhedral subdivision:
@@
    Let $C_{\min}$ be the minimal non-empty face of $C$ (i.e., its maximal affine subspace).
@@
    Define the height function: 0 on the vertices of $P$, and $1$ at $X$.
@@
Then for any $F\in\mathcal{P}_{B}$ we have $\operatorname{St}_{\mathcal{P}_{B}}(F)=\operatorname{St}_{\mathcal{X}}(\operatorname{Conv}(F,X))$. Thus, for any face $Q\in\mathcal{C}$ containing $F$, we have $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,F;t)=\ell_{\mathcal{P}}(\mathcal{X},Q,F;t)$.
Consider a face $U\in\mathcal{P}_{B}$ of $P$ and face $Q\in\mathcal{C}$ such that $U\subset Q$. Denote $\dim(U)=u$ and $\dim(Q)=q$.
@@
where $R_{<(q-u)/2}$ is a polynomial of degree less than (qu)/2(q-u)/2 whose coefficients are determined by the coefficients of $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$.
@@
The polynomial $\ell_{\mathcal{C}^{\pm}}(\mathcal{P},Q,U;t)$ has symmetric coefficients centered at (qu)/2(q-u)/2 and polynomial $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$ has symmetric coefficients centered at (qu1)/2(q-u-1)/2. Therefore, Lemma 3.11 fully recovers $\ell_{\mathcal{C}^{\pm}}(\mathcal{P},Q,U;t)$ from $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$. For instance, $\ell_{\mathcal{C}^{\pm}}(\mathcal{P},Q,U;t)$ has non-positive coefficients.
@@
    The first summand is the sum of products of pairs of $g$-polynomials of posets with rank sum qu+1q-u+1.
@@
Therefore, the first summand has degree less than (qu)/2(q-u)/2 and can be rewritten as $R_{<(q-u)/2}$.
The second summand is equal to $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$ and this completes the proof.
@@
Recall that by Remark 2.16, the local $h$-polynomial $\ell_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P},Q,U;t)$ has symmetric coefficients centered at (qu)/2(q-u)/2. Therefore, its coefficients are indeed determined by the lower-degree terms, which in turn are determined by $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$.
By Theorem 2.22 and Remark 3.10 the polynomial $\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},Q,U;t)$ has symmetric unimodal coefficients centered at (qu1)/2(q-u-1)/2. Hence, we obtained the following corollary.
@@
4 Generalization of $\mathbf{h^{*}}$- and $\mathbf{\ell^{*}}$-polynomial
4.1 $\mathbf{h^{*}}$- and $\mathbf{\ell^{*}}$-polynomial of polyhedral strong formal subdivision
An abstract lattice polyhedral complex $\mathcal{S}$ is a poset whose elements are ‘‘abstract’’ convex lattice polytopes, with the properties that: (1) every face of any polytope in $\mathcal{S}$ is also contained in $\mathcal{S}$, and (2) any two polytopes in $\mathcal{S}$ intersect along a common face. Formally, abstract lattice polyhedral complex is a map $\theta$ from a poset $\Gamma$ to lattice polytopes in (different) lattices such that for every $x\in\Gamma$ the lower set $\Gamma_{\leq x}$ is the face poset of the polytope $\theta(x)$ and for every $y\leq x$ there is a lattice isomorphism between the polytope $\theta(y)$ and the corresponding face of the polytope $\theta(x)$. We do not assume that there is a geometric realization in a common lattice.
@@ 4.1 $\mathbf{h^{*}}$- and $\mathbf{\ell^{*}}$-polynomial of polyhedral strong formal subdivision
It follows e.g. from [KS16, Theorem 3.11] that we recover the $h^{*}$-polynomial via the local $h^{*}$-polynomial:
@@
The first formula is implied by Definition 4.1 and the corresponding Katz-Stapledon decomposition formula from Lemma 2.33. The equivalence of these formulas is proved the same way as in [KS16, Lemma 7.12]. ∎
4.2 $\mathbf{h^{*}}$- and $\ell^{*}$-polynomial of agglutination